roulette wheel (see problem 5 of section 3.2 for a description of roulette) is spun twice. Find the probability of: (a) red first, black second (b) red and black. c) red or green first, red second. (d) the number 7 first, green second. (e) the number 4 and the number 8.
Answer #1

Solution:-

a) The probability that red first and black second is 0.2244

Number of red = 18

Number of black = 18

The probability that red first and black second = \frac{18}{38}\times \frac{18}{38}

b) The probability that red first and black second is 0.4488.

Number of red = 18

Number of black = 18

P(Red and black) = P(R, B) + P(B, R)

The probability that red and black =2\times \frac{18}{38}\times \frac{18}{38}

c) The probability that red or green first and black second is 0.2493

Number of red or green = 20

Number of red = 18

The probability that red or green first and red second = \frac{20}{38}\times \frac{18}{38}

d) The probability of number 7 first and green second is 0.00139.

Number of green = 2

The probability that number 7 first and green second = \frac{1}{38}\times \frac{2}{38}